the length of a rectangle is 6cm longer than its width if the perimeter of the rectangle is 52cm find its area

Answer by checkley79(3341)

L=W+6
PER=2L+2W
52=2(W+6)+2W
52=2W+12+2W
52-12=4W
40=4W
W=40/4
W=10 CM.
L=10+6
L=16 CM
PROOF:
52=2*10+2*16
52=20+32
52=52
AREA=L*W
AREA=10*16
AREA=160 CM^2